

********************************************************************
UARS WEEKLY STATUS REPORT - 21 May 2004
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
61% of each day's data amounting to about 541 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution,
MgII, occultation, solar pointing offset, and high resolution
continuous monitoring scans for periods of UARS real-time
telemetry or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1999119. MEPS data files between
1999120 and 1999159 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 131/1220 69292 OFF TO DAY-ONLY OPS
HALOE 133/1253 69323 POWER ON
SOLSTICE 134/1210 69337 DAY-ONLY TO FULLTIME
SUSIM 134/1533 69340 POWER ON
The beta angle ranged between the angles of 02.5 degrees to 24.5
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 36 degrees on 22 May 2004 (DOY 143).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 1.29 degrees C. Additional
battery performance data for each battery on 07 May and 14 May 2004
are provided in Appendix D.
The clock error ranged from 10.0 to -12.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
HALOE = 02
PEM = 01
SOLSTICE = 01
SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 06
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Six (6) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT began taking SSA only events on TDW on 23 April 2004. This
change to operations was needed due to the TDRS-8 move to TDW. Since
the SSA service is in higher demand than the MA service, the FOT
receives approximately 50% of the events requested. This effectively
yields the realtime support of 2.5 TDRS instead of 3 TDRS. The FOT
will continue to operate in this mode until the software fix to the
CMS is made to ingest the SMA view periods correctly. Preliminary
estimates project this fix to be released to operations no sooner than
the end of June.
MEETINGS
The FOT attended the HRDI attitude meeting on 12 May 2004 to hear the
HRDI teams findings and explore further possibilities for analysis.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 06 June 2004 (DOY 158). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 14 May 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
69251 129/1043 04-136 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
69279 130/1444 04-134 MPS YAW:PSU TO VT-L5
69280 130/1544 04-134 MPS YAW:BATSOC TO 100%
69292 131/1220 04-137 PEM OFF TO DAY-ONLYL OPS
69323 133/1253 04-138 HALOE POWER ON PART-1
69326 133/1830 04-140 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
69327 133/2012 04-139 HALOE POWER ON PART-2
69337 134/1210 04-141 SOLSTICE DAY-ONLY TO FULLTIME OPS
69340 134/1533 04-142 SUSIM POWER ON PART-1
69343 134/2124 04-143 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
69351 135/1026 04-142 SUSIM POWER ON (COMPLETE)
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 14 May 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 14 May 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
69251 129/1109 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
69326 133/1830 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
69343 134/2128 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 14 May 2004
07 May 2004 - GMT Day 128 Beta = 01.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +00.0/-184.8 +5.79/+4.50 +27.2/-13.6 25.0 1.00 21.6
3 +22.4/-072.8 +4.50/+3.23 +26.8/-14.4 25.0 1.00 21.3
14 May 2004 - GMT Day 135 Beta = 24.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-201.6 +5.79/+4.50 +28.0/-14.0 24.8 1.03 20.4
3 +11.2/-134.4 +4.50/+3.87 +27.2/-14.4 24.8 1.02 20.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
69279 130/1444 LV6 -> LV5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 14 May 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-024 FOT 69257 129/0323 30832 Generic Late Acquisition
04-024 FOT 69263 129/1305 30832 Generic Late Acquisition
04-025 FOT 69308 132/1357 35842 Data Dropout: data loss
04-026 FOT 69339 134/1458 35860 Data Dropouts
04-027 FOT 69344 134/2304 35864 Data Dropouts
04-024 FOT 69348 135/0447 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-024 FOT See List May 04 30832 Generic Late Acquisition
04-025 FOT 69308 132/1357 35842 Data Dropout: data loss
04-026 FOT 69339 134/1458 35860 Data Dropouts
04-027 FOT 69344 134/2304 35864 Data Dropouts